OpenStudy (anonymous):

A) if f(2) =4 can you conclude anything about the limit of f(x) as x approaches 2? explain your reasoning B) if the limit of f(x) as x approaches 2 is 4, can you conclude anything about f(2)? explain your reasoning

terenzreignz (terenzreignz):

What is your gut feel? lol

OpenStudy (anonymous):

idk

terenzreignz (terenzreignz):

For part (A) If \[\Large f(2) = 4\] Then does that mean \[\Large \lim_{x\rightarrow 2}f(x) = 4 \qquad \color{red}?\]

OpenStudy (anonymous):

not necessarily

terenzreignz (terenzreignz):

That's right :) Can you think of a counterexample? (HINT: Think piecewise...)

OpenStudy (anonymous):

nope....I'm having issues with piecewise functions as well

terenzreignz (terenzreignz):

Well, a most basic example would be \[\Large f(x) =\left\{\begin{matrix}4&&&x\le 2\\\\0&&&x>2\end{matrix}\right.\]

terenzreignz (terenzreignz):

Now clearly, f(2) = 4, but the limit doesn't exist as x approaches 2 because they approach different values from the left and from the right.

OpenStudy (anonymous):

can that be found if the question were graphed

terenzreignz (terenzreignz):

Possibly... IS the question graphed?

OpenStudy (anonymous):

no...but i learn better with visuals.... I'm a little slow

terenzreignz (terenzreignz):

Okay, what this piecewise function looks like on a graph is this: Say this is your Cartesian Plane

terenzreignz (terenzreignz):

 And it's clearer to see that f(2) = 4 but the function, as x approaches 2 from the LEFT, goes to 4, but when it approaches from the RIGHT, it goes to zero. thus, the limit doesn't exist :)

OpenStudy (anonymous):

ohhh i get it

terenzreignz (terenzreignz):

Okay? So part (a) is done? :)

OpenStudy (anonymous):

yea....you can say that

terenzreignz (terenzreignz):

I can say it, but we can't really move on until you can say it yourself :P Are you convinced about our conclusion with part (A) ?

OpenStudy (anonymous):

yea i can see it now

terenzreignz (terenzreignz):

Okay, part (B) then... the converse. Suppose we know that \[\Large \lim_{x\rightarrow 2}f(x) = 4\] does it mean that \[\Large f(2) = 4 \qquad \color{red}?\]

OpenStudy (anonymous):

I don't think so

terenzreignz (terenzreignz):

Can you think of a counterexample?

terenzreignz (terenzreignz):

You can think piecewise, but there are answers which I think are slightly prettier :3

OpenStudy (anonymous):

like what

terenzreignz (terenzreignz):

Well, first, maybe you want a piecewise function? :P

terenzreignz (terenzreignz):

Because the 'prettier' one is pretty hard to graph :P

terenzreignz (terenzreignz):

Here's the 'prettier' one (the one which isn't piecewise) \[\Large f(x) = \frac{x^2-4}{x-2}\]

terenzreignz (terenzreignz):

Clearly, f(2) won't exist, since the denominator would be made zero, right?

OpenStudy (anonymous):

right

terenzreignz (terenzreignz):

Now, you just have to show that the limit exists and that \[\Large \lim_{x\rightarrow2}\frac{x^2-4}{x-2}=4\] Can you do that? :)

terenzreignz (terenzreignz):

HINT: Numerator is factorable...

OpenStudy (anonymous):

yes... the top factors to become (x+2) (x-2) so the x-2's cancel and 2+2 equals 4

terenzreignz (terenzreignz):

Yup precisely :) So for this function \[\Large f(x)= \frac{x^2 -4}{x-2}\] The limit as x goes to 2 is indeed 4, but f(2) is not 4 (since it doesn't exist) That's it ^_^

terenzreignz (terenzreignz):

So actually, the answer to both of your questions is "no conclusion can be made" :)

OpenStudy (anonymous):

ohh ok...now i have another question.... how were you able to find the counterexample \[f \left( x \right)= \frac{ x^{2}-4 }{ x-2 }\]

terenzreignz (terenzreignz):

From experience :)

terenzreignz (terenzreignz):

That's like a textbook example showing that \[\Large \lim_{x\rightarrow a}f(x)\] is not necessarily equal to \[\Large f(a)\]

terenzreignz (terenzreignz):

I have to go now :) Hope you learned something today XD ------------------------------------------ Terence out

OpenStudy (anonymous):

ok thanks..this math lesson has been very helpful